OpenStudy (anonymous):

help :( Find the surface area and volume of a sphere. A = 4 r2 V = 4/3r3 A sphere has a radius of 4 inches.

OpenStudy (jdoe0001):

\(\bf \textit{Area of a sphere}=4\pi r^2\qquad \textit{Volume of a sphere}=\cfrac{4}{3}\pi r^3\\ \quad \\ r = radius\qquad \textit{in this case radius or}\quad r=4\quad thus\\ \quad \\ \textit{Area of a sphere}=4\pi (4)^2\qquad \textit{Volume of a sphere}=\cfrac{4}{3}\pi (4)^3\)